1.A truck and an ambulance move at the same rate of 13 m/s toward each other but on different lanes. If the ambulance siren emits sound of frequency 1200 Hz, what is the frequency of the sound that the truck driver hears?


2. The police car is moving toward a warehouse at 30 m/s. If the police car is emitting a 1kHz sound, what frequency does the driver hear reflected from the warehouse? Assume that the speed of sound is 340 m/s.


3. What is the frequency heard by a person driving at 15 m/s toward a factory whistle that is blowing at a frequency of 800 Hz? Assume that the speed of sound is 340 m/s


1
Expert's answer
2022-02-10T08:37:13-0500

The Doppler's effect equation says


"f=f_0\frac{c\pm v_r}{c\pm v_s}"

1


"f=\rm 1200\: Hz*\frac{340\: m/s+13\: m/s}{340\: m/s-13\: m/s}=1295\: Hz"


2


"f=\rm 1000\: Hz*\frac{340\: m/s}{340\: m/s-30\: m/s}=1097\: Hz"


3


"f=\rm 800\: Hz*\frac{340\: m/s+15\: m/s}{340\: m/s}=835\: Hz"